# Distinguishing Predictive and Generative AI in Regulation

### v2

@@ 4. Recommendations
A common critique of tech-neutral regulation is that it often fails to account for the distinct and significant consequences brought by technology (Hildner, 2006; Brenner, 2011). By the time the harms of a technology become apparent through widespread adoption, regulation may be ineffective and even impossible to implement (Gee, 2009). For instance, in nuclear safety regulation, the long-term environmental and health impacts from radioactive contamination and nuclear meltdowns are so severe and irreversible that a tech-neutral approach is inadequate (U.S. Nuclear Regulatory Commission, 2011). Instead, nuclear safety regulation relies on highly specific rules tailored to the unique risks of nuclear technology, such as strict access to nuclear materials and safety protocols that govern nuclear facilities (World Nuclear Association, 2021). Similarly, generative AI presents a unique challenge that predictive AI does not: risk of catastrophic harm (Hendrycks et al., 2023). Catastrophic harm refers to events with potentially devastating consequences for human society, such as the misuse of AI to create bioweapons or the loss of control over advanced AI systems (Sunstein, 2007). Unlike more predictable risks, catastrophic harms are difficult to quantify because they often involve scenarios with little to no historical data, making it nearly impossible to assign a precise probability of occurrence (Farber, 2011). Some caution that the race among companies to advance AI heightens the risk that these models could pose unforeseen threats, calling for pauses in AI development (Future of Life Institute, 2023). The concern that generative AI might be used for a broad range of dangerous and difficult-to-anticipate applications necessitates tech-specific regulation to preemptively mitigate these risks rather than reacting after the fact, when it may be too late to avert disaster.
@@
Second, tech-specific regulation should be reserved for cases where the potential harm is plausible and evidence-based, significant, and attributable to the specific technology or technological configuration. An example is AI-generated nonconsensual intimate imagery (NCII) based in the likeness of real people, otherwise known as deepfakes. AI-generated NCII is a current problem, with significant and well-documented harm (McGlynn et al., 2017; Henry et al., 2020; Laird et al., 2024). While NCII is starting to be addressed with criminalization, civil liability, and/or mandated takedowns (U.S. Congress, 2024), NCII may also be a candidate for tech-specific regulation because the harm is accomplished immediately, cheaply, and at scale, making it harder to detect and remove (Thiel et al., 2023; Internet Watch Foundation, 2024). An example of such tech-specific regulation would be a requirement that image generators have filters that prevent the generation of NCII with features of real people (National Institute of Standards and Technology, 2024). This is tech-specific because it regulates through a configuration of the technology, rather than addressing the harm directly, like liability for its creation or takedown requirements do (California State Legislature, 2024; Massachusetts General Court, 2024). So far, the technological limitations have not been totally successful—ability to evade them being one of the risks of tech-specific regulation—but something could be created that performs better, and successfully prevents the harm (Ding and Suresh, 2025).
@@
Third, tech-specific regulation may be justified when the risks posed by a specific technology are significant enough that the benefits of regulation interventions outweigh the costs. One example of this could be facial recognition technology (FRT). Scholars like Woodrow Hartzog and Evan Selinger have argued that FRT’s mere existence is uniquely dangerous in a way that justifies banning the technology outright, due in part to the unique role of the face in our identity and social interactions (Hartzog and Selinger, 2018). In fact, some state and local governments have already enacted bans on its use (Sheard and Schwartz, 2022). These restrictions curb the indiscriminate deployment of facial recognition technology by state actors, to prevent mass privacy violations, wrongful arrests, and disproportionate targeting of marginalized communities. While facial recognition technology may offer potential benefits in security and efficiency, ban advocates argue that these advantages are grossly overshadowed by the severe and evidenced risk it poses (Harwell and Dou, 2020; Hill, 2020). Generative AI may present analogous concerns in biosecurity. It risks facilitating a bio-engineered pandemic by (1) lowering the barrier of entry for non-experts to develop biological weapons, and (2) discovering novel, more lethal pathogens that evade current safeguards. The former may reduce the cost and expertise required to create harmful biological agents, while the latter raises the ceiling of harm by increasing the lethality and transmissibility of pathogens. These two outcomes establish distinct threat models and introduce different trade-offs that regulators will need to evaluate and prioritize when weighing the benefits of AI advancements for bioresearch against the costs of potential misuse.
@@
The current discourse on AI regulation has predominantly focused narrowly on the AI value chain. While this has yielded concrete technical governance proposals (OpenAI, 2024; Sastry et al., 2024; Browne, 2024), it neglects the systemic nature of AI risks and impacts, as its harm manifests through interactions across an ecosystem.
@@
One way to conceptualize these interactions is through a risk chain: a sequence of actions that traces how potential harms are conceived, developed, and realized across an ecosystem. In the case of biosecurity, a risk chain might begin with an actor having an “irresponsible, misguided, or malicious” intention. This intention gives rise to a biological idea, which is then converted into biological data—such as a pathogen genome—and transformed into a live biological artifact. The artifact is subsequently cultured, tested, and ultimately dispersed into the target environment (Sandberg and Nelson, 2020). In the following, we use bioweapons development as a case study to study how harm progresses or is prevented from one stage to another.
@@
Scholars have long observed that the particular arrangements and affordances of digital technology create extra-legal structures and contours that shape conduct in practice, often in ways that regulation and laws adapt to and build on (Reidenberg, 1998; Lessig, 2006; Surden, 2007; Ohm and Frankle, 2018). These structural ”constraints” function as “regulators of behavior that prevent conduct through technological or physical barriers in the world” (Surden, 2007). These constraints can be physical (e.g., access to infrastructure), economic (e.g., material costs), or epistemic (e.g., domain expertise). [^6] We note that these categories of barriers are not fixed or mutually exclusive. In practice, they often overlap and reinforce one another. For instance, a lack of technical knowledge (epistemic) may require hiring experts (economic), which in turn may be necessary to operate specialized systems (physical). While analytically distinct, these barriers often operate together as part of a broader regulatory environment. Much like law, these constraints can make certain activities prohibitively difficult or expensive to carry out at scale. They shape which harms can propagate along a risk chain and which are effectively contained by the ecosystem.
@@
Bioweapons development offer a clear example of how structural constraints operate within a risk chain. Even if an actors begins with a harmful biological idea, progressing through the chain requires access to biomaterials and specialized lab infrastructure (Peppin et al., 2025; Batalis, 2023). This requirement imposes physical constraints for obtaining and handling hazardous biomaterials and highlights risk chain dependencies, as the movement and use of biomaterials necessarily involve interactions with institutions that produce and distribute these agents (ASPR, 2025). These dependencies make certain transitions in the risk chain, such as transforming a digital genome into a cultured pathogen, contingent on navigating the physical, economic, and epistemic barriers. At the same time, these points offer opportunities for policy interventions, including screening and monitoring the distribution of biomaterials (Kane and Parker, 2024; ASPR, 2010), requiring licenses for researchers working with high-risk biological agents (Jonsson et al., 2013), and enforcing strict security protocols in laboratory facilities (National Research Council, 2009).
@@
We propose a three-step framework to help identify and evaluate how AI interacts with the risk chain. First, regulators should identify the key dependencies and constraints at each stage of the risk chain. These could include both deliberately engineered dependencies that make certain actions infeasible without authorization, as well as by-product constraints of the technological and physical environment raising the difficulty, cost, or time required to carry out harmful activities (Surden, 2007). In biosecurity, multiple constraints and dependencies impede the progression of the risk chain, making it more difficult to carry out biological attacks. There are epistemic barriers due to a lack of biological knowledge and scientific training (National Research Council, 2004; Ouagrham-Gormley, 2014). Next, actors need biomaterials, specialized equipement, state-of-the-art methods and deep domain expertise (Leitenberg et al., 2012). And the final stages of culturing, testing, and deploying biological agents are time- and resource-intensive (National Research Council, 2004; European Commission, 2005; Nuclear Threat Initiative, 2023)
@@
Second, regulators should assess the extent to which existing regulations depend on structural constraints to remain effective. When these constraints shift or weaken due to technological changes, the regulatory regimes built around them may lose their force. In the case of biosecurity, regulatory efforts to deter the creation of bioweapons have traditionally relied on physical constraints, such as controlling access to hazardous biomaterials and enforcing safety protocols within high-containment laboratory facilities (United Nations, nd; ASPR, 2025; NSTC, 2022). These efforts rest on the assumption that malicious actors cannot operationalize biological threats without overcoming significant physical barriers.
@@
The extent to which these constraints remain effective depends on the regulatory objectives and assumptions underlying the structure of the risk chain. For instance, even if generative AI lowers the barrier to entry to biodevelopment—by expanding access to biological information and reducing the time and cost needed for pathogen development—the physical constraints likely remain intact (Jefferson et al., 2014; Grushkin et al., 2013). Alternatively, if regulation focused on epistemic constraints, such as restricting access to sensitive methodological knowledge by redacting details in published biology manuscripts (Enserink, 2011), then the regulatory emphasis would shift to limiting knowledge diffusion. Regulators should determine which constraints are most critical to preserving the function of existing legal regimes and assess how AI may undercut their regulatory force.
@@ 5. Conclusion
Our goal in this work is to illustrate where the differences between generative and predictive AI are merely matters of degrees—where existing regulatory approaches remain applicable—and when they are differences in kind, demanding new regulatory instruments and forms of intervention. We identify four distinct areas of meaningful difference and make three recommendations for policymakers based on these differences. We believe that these recommendations will lead to more effective policymaking for generative AI and will avoid the fear-hype binary that seems to cast a shadow over AI policymaking today.

### v3

@@ 4. Recommendations
First, tech-specific regulation should only be used when tech-neutral regulation is impossible or impractical. A common critique of tech-neutral regulation is that it may be ineffective and even impossible to implement, by the time harms of a technology become apparent through widespread adoption (Gee, 2009). For instance, in nuclear safety regulation, the long-term environmental and health impacts from radioactive contamination and nuclear meltdowns are so severe and irreversible that a tech-neutral approach is inadequate (U.S. Nuclear Regulatory Commission, 2011). Instead, nuclear safety regulation relies on highly specific rules tailored to the unique risks of nuclear technology, such as strict access to nuclear materials and safety protocols that govern nuclear facilities (World Nuclear Association, 2021). Similarly, GPAI presents a severe consequence that task-specific AI does not: risk of catastrophic harm (Hendrycks et al., 2023). Catastrophic harm refers to events with potentially devastating consequences for human society, such as the misuse of AI to create bioweapons or the loss of control over advanced AI systems (Sunstein, 2007). Unlike more predictable risks, catastrophic harms are difficult to quantify because they often involve scenarios with little to no historical data, making it nearly impossible to assign a precise probability of occurrence (Farber, 2011). Some caution that the race among companies to advance AI heightens the risk that these models could pose unforeseen threats, calling for pauses in AI development (Future of Life Institute, 2023). The concern that GPAI might be used for a broad range of dangerous and difficult-to-anticipate applications necessitates tech-specific regulation to preemptively mitigate these risks rather than reacting after the fact, when it may be too late to avert disaster.
@@
Second, tech-specific regulation should be reserved for cases where the potential harm is plausible and evidence-based, significant, and attributable to the specific technology or technological configuration. An example is AI-generated nonconsensual intimate imagery (NCII) based in the likeness of real people, otherwise known as deepfakes. AI-generated NCII is a current problem, with significant and well-documented harm (McGlynn et al., 2017; Henry et al., 2020; Laird et al., 2024). While NCII is starting to be addressed with criminalization, civil liability, and/or mandated takedowns (U.S. Congress, 2024), NCII may also be a candidate for tech-specific regulation because the harm is accomplished immediately, cheaply, and at scale, making it harder to detect and remove (Thiel and others, 2023; Internet Watch Foundation, 2024). An example of such tech-specific regulation would be a requirement that image generators have filters that prevent the generation of NCII with features of real people (National Institute of Standards and Technology, 2024). This is tech-specific because it regulates through a configuration of the technology, rather than addressing the harm directly, like liability for its creation or takedown requirements do (California State Legislature, 2024; Massachusetts General Court, 2024). So far, the technological limitations have not been totally successful—ability to evade them being one of the risks of tech-specific regulation—but something could be created that performs better, and successfully prevents the harm (Ding and Suresh, 2025).
@@
Third, tech-specific regulation may be justified when the risks posed by a specific technology are significant enough that the benefits of regulation interventions outweigh the costs. One example of this could be facial recognition technology (FRT). Hartzog and Selinger have argued that FRT’s mere existence is uniquely dangerous in a way that justifies banning the technology outright, due in part to the unique role of the face in our identity and social interactions (Hartzog and Selinger, 2018). In fact, some state and local governments have already enacted bans on its use (Sheard and Schwartz, 2022). These restrictions curb the indiscriminate deployment of facial recognition technology by state actors, to prevent mass privacy violations, wrongful arrests, and disproportionate targeting of marginalized communities. While facial recognition technology may offer potential benefits in security and efficiency, ban advocates argue that these advantages are grossly overshadowed by the severe and evidenced risk it poses (Harwell and Dou, 2020; Hill, 2020). GPAI may present analogous concerns in biosecurity. It risks facilitating a bio-engineered pandemic by (1) lowering the barrier to entry for non-experts to develop biological weapons, and (2) discovering novel, more lethal pathogens that evade current safeguards. The former may reduce the cost and expertise required to create harmful biological agents, while the latter raises the ceiling of harm by increasing the lethality and transmissibility of pathogens. These two outcomes establish distinct threat models and introduce different trade-offs that regulators will need to evaluate and prioritize when weighing the benefits of AI advancements for bioresearch against the costs of potential misuse.
@@
The current discourse on AI regulation has predominantly focused narrowly on the AI value chain. While this has yielded concrete technical governance proposals (OpenAI, 2024a; Sastry et al., 2024; Browne, 2024), it neglects the systemic nature of AI risks and impacts, as its harm manifests through interactions across an even broader ecosystem. We propose an evaluation framework to assess regulatory approaches across domains, using biosecurity as a case study. The framework examines where existing regulations assume conditions that GPAI now disrupt, and pinpoints where new measures are needed to close new gaps.
@@
One way to conceptualize the interactions between actors, systems, and resources across an ecosystem is through a risk chain: a sequence of actions that traces how potential harms are conceived, developed, and realized across an ecosystem. In biosecurity, a risk chain might begin with an actor having an “irresponsible, misguided, or malicious” intention (Sandberg and Nelson, 2020). This intention gives rise to a biological idea, which is then converted into biological data—such as a pathogen genome—and transformed into a live biological artifact. The artifact is subsequently cultured, tested, and ultimately dispersed into the target environment (Sandberg and Nelson, 2020). In the following, we use bioweapons development as a case study to study how harm progresses or is prevented from one stage to another.
@@
Scholars have long observed that the particular arrangements and affordances of digital technology create extra-legal structures and contours that shape conduct in practice, often in ways that regulation and laws adapt to and build on (Reidenberg, 1998; Lessig, 2006; Surden, 2007; Ohm and Frankle, 2018). These structural “constraints” function as “regulators of behavior that prevent conduct through technological or physical barriers in the world” (Surden, 2007). These constraints can be physical (e.g., access to infrastructure), economic (e.g., material costs), or epistemic (e.g., domain expertise). [^7] We note that these categories of barriers are not fixed or mutually exclusive. In practice, they often overlap and reinforce one another. For instance, a lack of technical knowledge (epistemic) may require hiring experts (economic), which in turn may be necessary to operate specialized systems (physical). While analytically distinct, these barriers often operate together as part of a broader regulatory environment. Much like law, these constraints can make certain activities prohibitively difficult or expensive to carry out at scale. They shape which harms can propagate along a risk chain and which are effectively contained by the ecosystem.
@@
Bioweapons development offer a clear example of how structural constraints operate within a risk chain. Even if an actor begins with a harmful biological idea, progressing through the chain requires access to biomaterials and specialized lab infrastructure (Peppin et al., 2025; Batalis, 2023). This requirement imposes physical constraints for obtaining and handling hazardous biomaterials and highlights risk chain dependencies, as the movement and use of biomaterials necessarily involve interactions with institutions that produce and distribute these agents (ASPR, 2025). These dependencies make certain transitions in the risk chain, such as transforming a digital genome into a cultured pathogen, contingent on navigating the physical, economic, and epistemic barriers. At the same time, these points offer opportunities for policy interventions, including screening and monitoring the distribution of biomaterials (Kane and Parker, 2024; ASPR, 2010), requiring licenses for researchers working with high-risk biological agents (Jonsson et al., 2013), and enforcing strict security protocols in laboratory facilities (National Research Council, 2009).
@@
We propose a three-step domain-independent framework to help identify and evaluate how AI interacts with the risk chain. First, regulators should identify the key dependencies and constraints at each stage of the risk chain. These could include both deliberately engineered dependencies that make certain actions infeasible without authorization, as well as by-product constraints of the technological and physical environment raising the difficulty, cost, or time required to carry out harmful activities (Surden, 2007). In biosecurity, multiple constraints and dependencies impede the progression of the risk chain, making it more difficult to carry out biological attacks. To being, there are epistemic barriers due to a lack of biological knowledge and scientific training (National Research Council, 2004; Ouagrham-Gormley, 2014). Next, actors need biomaterials, specialized equipement, state-of-the-art methods and deep domain expertise (Leitenberg et al., 2012). And the final stages of culturing, testing, and deploying biological agents are time- and resource-intensive (National Research Council, 2004; European Commission, 2005; Nuclear Threat Initiative, 2023)
@@
Second, regulators should assess the extent to which existing regulations depend on structural constraints to remain effective. When these constraints shift or weaken due to technological changes, the regulatory regimes built around them may lose their force. In the case of biosecurity, regulatory efforts to deter the creation of bioweapons have traditionally relied on physical constraints, such as controlling access to hazardous biomaterials and enforcing safety protocols within high-containment laboratory facilities (United Nations, n.d.; ASPR, 2025; NSTC, 2022). These efforts rest on the assumption that malicious actors cannot operationalize biological threats without overcoming significant physical barriers.
@@
The extent to which these constraints remain effective depends on the regulatory objectives and assumptions underlying the structure of the risk chain. For instance, even if GPAI lowers the barrier to entry to biodevelopment—by expanding access to biological information and reducing the time and cost needed for pathogen development—the physical constraints likely remain intact (Jefferson et al., 2014; Grushkin et al., 2013). Alternatively, if regulation focused on epistemic constraints, such as restricting access to sensitive methodological knowledge by redacting details in published biology manuscripts (Enserink, 2011), then the regulatory emphasis would shift to limiting knowledge diffusion. Regulators should determine which constraints are most critical to preserving the function of existing legal regimes and assess how AI may undercut their regulatory force.
@@ 5. Conclusion
Our goal in this work is to illustrate when the differences between general-purpose and task-specific AI are merely matters of degrees—where existing regulatory approaches remain applicable—and when they are differences in kind, demanding new regulatory instruments and forms of intervention. We identify four distinct areas of meaningful difference and make three recommendations for policymakers based on these differences. We believe that these recommendations can help policymakers avoid many of the fault lines in current AI policy debates and lead to more effective policymaking for GPAI.